jenna has a total of 23 coins, made up of dimes and nickels. if she has a $1.90, how many nickels and dimes does she have?
Answer by stanbon(75887)

jenna has a total of 23 coins, made up of dimes and nickels. if she has a $1.90, how many nickels and dimes does she have?
-----
Equations:
d + n = 23 coins
10d+5n=190 cents
-----------------------
Divide thru the lower equation by 5:
------
d + n = 23
2d + n = 38
----
Subtract and solve for "d":
d = 15 (# of dimes)
----
Solve for "n":
---
d + n = 23
15 + n = 23
n = 8 (# of nickels)
------------------------
Cheers,
Stan H.
=================